OpenStudy (anonymous):

Write the function in slope-intercept form: -5x + y = 7

OpenStudy (nincompoop):

slope intercept form looks like this y = mx+b

OpenStudy (nurali):

-5x + y = 7 y=mx+b y=7+5x y=5x+7 m=slope m=5

OpenStudy (anonymous):

you add 5 x and you get y = 5x + 7

OpenStudy (nincompoop):

since when are we allowed to give out answers? it doesn't matter if you gave out the solution! work on a different type of problem so that way you CAN KNOW if they understood it or not before moving to another question.

OpenStudy (anonymous):

Thanks you guys.. i was making it harder than it really is.

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

nincompoop it helps them understand and if you show the steps it helps them even more